# Scaling LLM Planning: NL2FLow for Parametric Problem Generation and Rigorous Evaluation

### v5

@@ 2 Method
This study used an automated planning integration system, NL2Flow, to generate a dataset of workflow planning problems and evaluate the performance of large language models in generating workflows and translating workflow problems in natural language. The NL2Flow’s profiler module (profiler.generators.batch_data_generator) was used to programmatically generate a diverse set of planning scenarios, varying parameters. This allowed for controlled generation of data suitable for evaluating LLM performance across a range of workflow configurations. LLM inference was conducted on compute clusters equipped with NVIDIA H100, H200, and A100 GPUs (with larger models, for example, DeepSeek-V3, using up to 8 H200 GPUs with 400GB memory). LLM-generated workflows were evaluated using NL2Flow’s Debugger module (nl2flow.debug.debug.BasicDebugger), which assesses soundness, validity, and optimality. A description of the AI planning principles is provided in the Appendix.
@@ 2.1 Domain Description
Automated workflow completion was a core requirement for the intelligent systems investigated in this study. I focused on the NL2Flow domain, which combined service composition with goal-oriented dialogue to generate task workflows (chakraborti2022natural, ). The objective was to develop a system capable of autonomously selecting and composing a sequence of actions, which are termed ‘custom’ actions, representing available services, to achieve user goals (e.g., submitting an expense report) (Listing 1). Each ‘custom‘ action was modeled as a parameterized function, and successful execution necessitated providing values for a defined set of input parameters. Critically, these parameter values were frequently unknown at the initiation of a workflow, representing information the system needed to acquire from the user. To address this, the system used a dialogue-based approach. When a parameter value was missing, the system generated an ‘ask’ action, a dialogue act specifically designed to elicit the necessary information. This process of identifying and requesting missing parameter values was analogous to slot-filling, a common technique in task-oriented dialogue systems, and ensured the system possessed the information required to progress. Successfully executing a ‘custom‘ action updated the system’s internal state, transitioning the corresponding parameter from unknown to known, and enabling the system to proceed with subsequent actions in the workflow.
@@ 2.2 Parametric Data Generation
A parametric workflow problem generator was developed to create an intermediate representation (IR) of the problem space. Implemented in Python, this generator allows for controlled variation in workflow complexity and characteristics. A dataset of 2,296 workflows was generated using a fixed random seed (20250429) to ensure reproducibility. Workflow generation was governed by six key parameters (Table 1): the number of ‘custom‘ actions (representing available operations), ‘custom‘ action arity (the number of input/output parameters defining each action’s interface), the number of goal actions (specifying the desired outcome), the proportion of slot-fillable variables (variables requiring values during planning), the complexity of variable mapping, and the proportion of coupled actions (actions where outputs serve as inputs to subsequent actions, creating dependencies). I aimed to generate ten unique samples for each valid parameter combination. ‘custom‘ actions were defined as callable components with specified input and output parameter sets, invoked when all required input parameters are known, and producing corresponding output parameters. These actions represent operations commonly found in agentic frameworks, such as agents and tools.
@@
To construct a workflow problem, 30 variables were generated and designated as either slot-fillable or initially known. ‘custom‘ actions were then created and connected to establish dependencies, with action outputs serving as inputs for subsequent actions. Variables not directly used within any action were assigned pre-defined values. A ‘custom‘ action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the ‘custom‘ action set. To focus on core workflow generation challenges, we excluded variable mapping and initial variable values from this study. The resulting intermediate representation (Python object) was compiled into PDDL strings using NL2Flow, enabling symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication.
@@
  • available_agentsavailable\_agents represent the available ‘custom‘ actions
@@ 2.3 Plan Evaluation
To assess the soundness of generated plans, the pddl compilation omits goal achievement. NL2Flow’s approach distinguishes between soundness and validity. For soundness checking, the planning problem was compiled into PDDL, but instead of including the original goals as requirements, the actions reported as executed within the LLM’s response were incorporated. This approach tests the LLM’s claimed actions actually executable. Validity was checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan. In this low-difficulty workflow problem setting, a plan was invalid, also, if the system prompted the user for a non-slot-fillable variable, or if a ‘custom‘ action was invoked with unknown input parameter values. I investigated the use of VAL (howey2004val, ) for plan validation; however, VAL did not support the PDDL language features necessary to represent the complexities introduced by NL2Flow’s PDDL compilation process, precluding its use in this evaluation.
@@ 3.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 3) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid plans was largely attributable to its ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of ‘custom‘ actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – showed that factors influencing the LLM’s failure to correctly identify ’no plan’ scenarios varied depending on the model and the prompt style (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort in Appendix). However, no single variable consistently predicted failure across all models or prompt styles.
@@ 3.1.4 Feasible Plan
When restricted to problems with feasible plans, Llama-3.1-405B-instruct-fp8 demonstrated the strongest performance, achieving the high rates of generating sound (91%), valid (85%), and optimal plans (69%) using the concise prompt style (Figure 4). I further investigated the influence of problem characteristics – including optimal plan length, number of ‘custom‘ actions, arity, coupling, proportion of slot-fillable variables, and number of goals – on the generation of valid plans. The relationship between these explanatory variables and plan quality (soundness, validity, and optimality) varied considerably across language models.
@@ 5 Conclusion
For the planning task, I investigated the impact of prompt style, comparing ‘verbose‘ (Listing A5) and ‘concise‘ (Listing A6) prompt styles. The ‘verbose‘ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ‘concise‘ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. Listings A3 and A4 show the description of ‘custom‘ action a__1a\_\_1 in verbose and concise prompts, respectively. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt.
Listing A3: ‘custom‘ action in verbose prompt
@@
Listing A4: ‘custom‘ action in concise prompt
@@
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as ’noplan’
@@
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as ’noplan’
@@
For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. The initial intermediate representation schema used nested structures to represent ‘custom‘ actions; however, this complexity proved unnecessary, as the same information could be represented with a simplified schema. To evaluate the capacity of LLMs to translate natural language descriptions of workflow problems into JSON, I used a streamlined schema (Listing A8) defined as follows:
  • actionsactions represents ‘custom‘ actions.
  • mappingsmappings corresponds to the ’mappings’ present in the intermediate representation.
@@
  • goal_action_idsgoal\_action\_ids corresponds to the ’goal_action_ids’ in the intermediate representation.
An example of this schema, used for evaluating LLM translation capacity, is provided as ‘translation_object.json‘ in the supplementary material.
@@
Many constants are defined for each type. The ’generic’ constants (v__xv\_\_x) represent specific instances of the objects being processed. The hasdonestatehas{-}done{-}state constants track the temporal status of actions. The constraintstatusconstraint{-}status and datumstatedatum{-}state constants define the possible values for these attributes. operatoroperator constants (a__xa\_\_x) represent specific, pre-defined actions.
@@
The domain defines two action types. ‘custom‘ actions are named according to the pattern a__xa\_\_x.
@@
Listing A9: ‘custom‘ action example in PDDL
@@
For example, Listing A9 illustrate the structure of ‘custom‘ actions defined in the domain model. This action, a__2a\_\_2, has no parameters. Its precondition requires that the action hasn’t been performed in the past (not(has_donea__2past)not(has\_done\>a\_\_2\>past)) and that a variable v__7v\_\_7 is known with certainty (knownv__7certainknown\>v\_\_7\>certain). The action’s effect marks the action as done in the present (has_donea__2presenthas\_done\>a\_\_2\>present), indicates v__7v\_\_7 has been used, makes v__9v\_\_9 free, establishes v__9v\_\_9 as known with certainty, indicates v__9v\_\_9 is no longer mapped, and increases the $total\mbox{-}cost$ by 10.
@@
Listing A10: ‘ask‘ action example in PDDL
@@
The initial condition of the problem, as described in ‘problem.pddl‘ in supplementary material, can be summarized as follows:
@@
  • For any generic object v__xv\_\_x used as an input parameter to a ‘custom‘ action, the predicate $(been_{u}sed\>v\_\_x)$ is true.
@@
In this study, the initial state was defined by a set of variables (v__xv\_\_x), each assigned a slot_goodnessslot\_goodness value of 1000000 if slot-fillable, and 150000 if not. These differing slot_goodnessslot\_goodness values represented a ’soft’ constraint, which was not used in this study because the precondition for the ‘ask‘ action prevented the system from requesting information for variables that were not slot-fillable. The mapped_tomapped\_to predicate establishes a self-mapping for each resource, essentially stating that each resource is initially assigned to itself. A new_itemnew\_item predicate indicates the creation of a new resource. The goal is to achieve conditions to complete a task or tasks (and(has_donea__xpresent)(has_donea__xpresent)and\>(has\_done\>a\_\_x\>present)(has\_done\>a\_\_x\>present)) while minimizing the totalcosttotal-cost.

### v6

@@ 2 Method
This study used an automated planning integration system, NL2Flow, to generate a dataset of workflow planning problems and evaluate the performance of large language models in generating workflows and translating workflow problems in natural language. The NL2Flow’s profiler module (profiler.generators.batch_data_generator) was used to programmatically generate a diverse set of planning scenarios, varying parameters. This allowed for controlled generation of data suitable for evaluating LLM performance across a range of workflow configurations. The code for running experiments is available in the NL2FLOW-Runner repository (https://github.com/IBM/nl2flow-runner). LLM inference was conducted on compute clusters equipped with NVIDIA H100, H200, and A100 GPUs (with larger models, for example, DeepSeek-V3, using up to 8 H200 GPUs with 400GB memory). LLM-generated workflows were evaluated using NL2Flow’s Debugger module (nl2flow.debug.debug.BasicDebugger), which assesses soundness, validity, and optimality. A description of the AI planning principles is provided in the Appendix.
@@ 2.1 Domain Description
Automated workflow completion was a core requirement for the intelligent systems investigated in this study. I focused on the NL2Flow domain, which combined service composition with goal-oriented dialogue to generate task workflows (chakraborti2022natural, ). The objective was to develop a system capable of autonomously selecting and composing a sequence of actions, which are termed ‘custom’ actions, representing available services, to achieve user goals (e.g., submitting an expense report) (Listing 1). Each ‘custom’ action was modeled as a parameterized function, and successful execution necessitated providing values for a defined set of input parameters. Critically, these parameter values were frequently unknown at the initiation of a workflow, representing information the system needed to acquire from the user. To address this, the system used a dialogue-based approach. When a parameter value was missing, the system generated an ‘ask’ action, a dialogue act specifically designed to elicit the necessary information. This process of identifying and requesting missing parameter values was analogous to slot-filling, a common technique in task-oriented dialogue systems, and ensured the system possessed the information required to progress. Successfully executing a ‘custom’ action updated the system’s internal state, transitioning the corresponding parameter from unknown to known, and enabling the system to proceed with subsequent actions in the workflow.
@@ 2.2 Parametric Data Generation
A parametric workflow problem generator was developed to create an intermediate representation (IR) of the problem space. Implemented in Python, this generator allows for controlled variation in workflow complexity and characteristics. A dataset of 2,296 workflows was generated using a fixed random seed (20250429) to ensure reproducibility. Workflow generation was governed by six key parameters (Table 1): the number of ‘custom’ actions (representing available operations), ‘custom’ action arity (the number of input/output parameters defining each action’s interface), the number of goal actions (specifying the desired outcome), the proportion of slot-fillable variables (variables requiring values during planning), the complexity of variable mapping, and the proportion of coupled actions (actions where outputs serve as inputs to subsequent actions, creating dependencies). I aimed to generate ten unique samples for each valid parameter combination. ‘custom’ actions were defined as callable components with specified input and output parameter sets, invoked when all required input parameters are known, and producing corresponding output parameters. These actions represent operations commonly found in agentic frameworks, such as agents and tools.
@@
To construct a workflow problem, 30 variables were generated and designated as either slot-fillable or initially known. ‘custom’ actions were then created and connected to establish dependencies, with action outputs serving as inputs for subsequent actions. Variables not directly used within any action were assigned pre-defined values. A ‘custom’ action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the ‘custom’ action set. To focus on core workflow generation challenges, we excluded variable mapping and initial variable values from this study. The resulting intermediate representation (Python object) was compiled into PDDL strings using NL2Flow, enabling symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication.
@@
  • available_agentsavailable\_agents represent the available ‘custom’ actions
@@ 2.3 Plan Evaluation
To assess the soundness of generated plans, the pddl compilation omits goal achievement. NL2Flow’s approach distinguishes between soundness and validity. For soundness checking, the planning problem was compiled into PDDL, but instead of including the original goals as requirements, the actions reported as executed within the LLM’s response were incorporated. This approach tests the LLM’s claimed actions actually executable. Validity was checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan. In this low-difficulty workflow problem setting, a plan was invalid, also, if the system prompted the user for a non-slot-fillable variable, or if a ‘custom’ action was invoked with unknown input parameter values. I investigated the use of VAL (howey2004val, ) for plan validation; however, VAL did not support the PDDL language features necessary to represent the complexities introduced by NL2Flow’s PDDL compilation process, precluding its use in this evaluation.
@@ 3.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 3) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid plans was largely attributable to its ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of ‘custom’ actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – showed that factors influencing the LLM’s failure to correctly identify ’no plan’ scenarios varied depending on the model and the prompt style (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort in Appendix). However, no single variable consistently predicted failure across all models or prompt styles.
@@ 3.1.4 Feasible Plan
When restricted to problems with feasible plans, Llama-3.1-405B-instruct-fp8 demonstrated the strongest performance, achieving the high rates of generating sound (91%), valid (85%), and optimal plans (69%) using the concise prompt style (Figure 4). I further investigated the influence of problem characteristics – including optimal plan length, number of ‘custom’ actions, arity, coupling, proportion of slot-fillable variables, and number of goals – on the generation of valid plans. The relationship between these explanatory variables and plan quality (soundness, validity, and optimality) varied considerably across language models.
@@ 5 Conclusion
6 Acknowledgements
The author gratefully acknowledges the close collaboration with Dr. Tathagata Chakraborti, whose guidance was instrumental in the design and implementation of NL2FLOW, particularly in the development of its core planning algorithm. The author also thanks Dr. Junkyu Lee at IBM Research for providing valuable feedback on an early draft of this manuscript.
@@ 6 Acknowledgements
For the planning task, I investigated the impact of prompt style, comparing ‘verbose’ (Listing A5) and ‘concise’ (Listing A6) prompt styles. The ‘verbose’ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ‘concise’ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. Listings A3 and A4 show the description of ‘custom’ action a__1a\_\_1 in verbose and concise prompts, respectively. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt.
Listing A3: ‘custom’ action in verbose prompt
@@
Listing A4: ‘custom’ action in concise prompt
@@
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as no plan
@@
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as no plan
@@
For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. The initial intermediate representation schema used nested structures to represent ‘custom’ actions; however, this complexity proved unnecessary, as the same information could be represented with a simplified schema. To evaluate the capacity of LLMs to translate natural language descriptions of workflow problems into JSON, I used a streamlined schema (Listing A8) defined as follows:
  • actionsactions represents ‘custom’ actions.
  • mappingsmappings corresponds to the ‘mappings’ present in the intermediate representation.
@@
  • goal_action_idsgoal\_action\_ids corresponds to the ‘goal_action_ids’ in the intermediate representation.
An example of this schema, used for evaluating LLM translation capacity, is provided as ‘translation_object.json’ in the supplementary material.
@@
Many constants are defined for each type. The ‘generic’ constants (v__xv\_\_x) represent specific instances of the objects being processed. The hasdonestatehas{-}done{-}state constants track the temporal status of actions. The constraintstatusconstraint{-}status and datumstatedatum{-}state constants define the possible values for these attributes. operatoroperator constants (a__xa\_\_x) represent specific, pre-defined actions.
@@
The domain defines two action types. ‘custom’ actions are named according to the pattern a__xa\_\_x.
@@
Listing A9: ‘custom’ action example in PDDL
@@
For example, Listing A9 illustrate the structure of ‘custom’ actions defined in the domain model. This action, a__2a\_\_2, has no parameters. Its precondition requires that the action hasn’t been performed in the past (not(has_donea__2past)not(has\_done\>a\_\_2\>past)) and that a variable v__7v\_\_7 is known with certainty (knownv__7certainknown\>v\_\_7\>certain). The action’s effect marks the action as done in the present (has_donea__2presenthas\_done\>a\_\_2\>present), indicates v__7v\_\_7 has been used, makes v__9v\_\_9 free, establishes v__9v\_\_9 as known with certainty, indicates v__9v\_\_9 is no longer mapped, and increases the $total\mbox{-}cost$ by 10.
@@
Listing A10: ‘ask’ action example in PDDL
@@
The initial condition of the problem, as described in ‘problem.pddl’ in supplementary material, can be summarized as follows:
@@
  • For any generic object v__xv\_\_x used as an input parameter to a ‘custom’ action, the predicate $(been_{u}sed\>v\_\_x)$ is true.
@@
In this study, the initial state was defined by a set of variables (v__xv\_\_x), each assigned a slot_goodnessslot\_goodness value of 1000000 if slot-fillable, and 150000 if not. These differing slot_goodnessslot\_goodness values represented a ‘soft’ constraint, which was not used in this study because the precondition for the ‘ask’ action prevented the system from requesting information for variables that were not slot-fillable. The mapped_tomapped\_to predicate establishes a self-mapping for each resource, essentially stating that each resource is initially assigned to itself. A new_itemnew\_item predicate indicates the creation of a new resource. The goal is to achieve conditions to complete a task or tasks (and(has_donea__xpresent)(has_donea__xpresent)and\>(has\_done\>a\_\_x\>present)(has\_done\>a\_\_x\>present)) while minimizing the totalcosttotal-cost.